Shift sum. Day 10/3/99

Richard Ford (rickford@fnal.gov)
Sun, 3 Oct 1999 15:47:49 -0500 (CDT)

Shift Summary for day shift, 03/Oct/99

On Shift: Theo, RickF + RickK + JasonH

Hours of beam delivered at all intensities: 8.0
Hours of beam delivered at nominal intensity: 7.3
Hours of E799 data (special runs not included): 7.0
Hours of special runs ( muon runs...):

Accounting of lost time: Start new run when l3/cpu rates were off

========================================================================
Runs ended this shift:

Run 14795 - Beam 799 - 25 M events
Run 14796 - Beam muon
Run 14797 - Beam 799 - ongoing

Tapes brought to FCC this shift:
(including tapes for online split)
UPA-C 058-60
UPS020

========================================================================
Accesses during shift:none

========================================================================
Overview of shift: Smooth running for the most part. 2 Hz spill quality
has been fair at best. Getting random scaler/sem errors.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14797 ??? 07:00 Inherit run 14797 already in progress

284 09:15 L3 statistics vs CPU indicates a problem with CPU
Stop run 14797. Fresh_start

14798 09:28 Start run 14798. Updated ADGN constants

??? 11:00 Noticed damp window is hanging. Trying to find expert.

12:30 BobT calls. Says raise intensity to 9e12. Done

12:40 JH walks in with the rosetta stone of damp restarting
Just do: $KTEVANA_DIR/bin/start_damp from
da_ktev3 account

186 13:35 Beam disappears. Problems with TRF2 water heater

195 13:50 Beam returns but at lower intensity
MCR Elog: We have bypassed the abort input for TRF2
and are attempting to run at 2 turns

14:30 Beam back to nominal.

15:30 End with 14798 still running
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

EXPERTS NOTE:
Noticed when going thru online kumacs that RC6 chan 7 shows frowny face
(high ADC). A couple MU3X,Y channels want their HV lowered. (frowny face again)

General Note: Getting scaler errors/sem (low limit) randomly thruout the
shift. Plotted sem vs different intensity monitors; nothing seems to have
changed in the last few days. Possibly an intermittant tgting problem.
BobT notified.